Shown here is an income statement in the traditional format for a firm with a sales volume of 18,000 units:

Revenues $ 144,000
Cost of goods sold ($9,000 + $2.80/unit) 59,400
Gross profit $ 84,600
Operating expenses:
Selling ($2,250 + $0.85/unit) 17,550
Administration ($4,900 + $0.35/unit) 11,200
Operating income $ 55,850


Required:

  1. Prepare an income statement in the contribution margin format.
  2. Calculate the contribution margin per unit and the contribution margin ratio.
  3. Calculate the firm's operating income (or loss) if the volume changed to
  1. 23,000 units.
  2. 10,000 units.
  1. Refer to your answer to part a when total revenues were $144,000. Calculate the firm's operating income (or loss) if unit selling price and variable expense per unit do not change and total revenues
  1. Increase by $13,500.
  2. Decrease by $7,500.
Answer #1
Contribution format Income statement
Sales             144,000
Less: Variable costs
Cost of goods sold               50,400
Selling expenses               15,300
Admin expenses                 6,300
Total variable costs               72,000
Contribution Margin               72,000
Less: Fixed costs
Cost of goods sold                 9,000
Selling expenses                 2,250
Admin expenses                 4,900
Total fixed costs               16,150
Net operating income               55,850
Contribution Margin per unit                         4
CM Ratio 50%
Operating income at 23000 units = 23000*4 - 16150 = $75,850
Operating income at 10000 units = 10000*4 - 16150 = $23,850
Operating income when revenue increases by $13500 = (144000+13500)*50% - 16150 = $62,600
Operating income when revenue decreases by $7500 = (144000-7500)*50% - 16150 = $52,100

Α B 144000 = 2.8*18000 =0.85*18000 =0.35*18000 =SUM(B4:36) =B2-B7 1 Contribution format Income statement 2 Sales 3 Less: Vari